On June 1, 20--, a depreciable asset was acquired for $4,020. The asset has an estimated useful life of five years (60 months) and no salvage value.

Using the straight-line depreciation method, calculate the book value as of December 31, 20--. If necessary, round your answer to two decimal places.
Answer #1

Answer

>Book Value = Cost of Asset - Accumulated Depreciation

>Working

A Cost $4,020
B Salvage Value $0
C = A - B Depreciable base $4,020
D Useful lufe in months 60
E = C/D Depreciation expense per month $67
F No. of month from June 1 to Dec 31 7
G = E x F Depreciation expense for 7 months $469
H = A - G Book Value on 31 Dec $3,551 = Answer

>Correct Answer: Book Value = $ 3,551